### v1

# Non-affine Families of 8 x 8 Complex Hadamard Matrices

Abstract
Six non-affine 3-parameter families of complex Hadamard matrices of order 8 are presented. These families contain Hadamard matrices that are not equivalent to any previously known Hadamard matrices in the literature. Each family arises from unimodular points of an affine variety defined by palindromic polynomials. The families are given as an image of a function that solves the corresponding system of polynomials on a domain that guarantees unimodularity of the solutions.
1 Introduction
Let $\mathbb{T}$ denote the set of unimodular complex numbers, that is, complex numbers with modulus 1. Let $M_{n}(\mathbb{T})$ be the set of $n\times n$ matrices over $\mathbb{T}$. A matrix $H\in M_{n}(\mathbb{T})$ is called a complex Hadamard matrix of order $n$ if it satisfies the equation $HH^{\dagger}=nI$, where $H^{\dagger}$ is the conjugate transpose of $H$ and $I$ is the identity matrix. This is equivalent to the condition that the rows (and columns) of $H$ are orthogonal in $\mathbb{C}^{n}$. The set of complex Hadamard matrices of order $n$ is denoted by $\mathbb{H}(n).$
A matrix $B\in\mathbb{H}(n)$ whose entries are $q$th roots of unity is said to be of Butson-type, abbreviated as BH$(n,q)$. For every order $n$, the Fourier matrix $F_{n}:=[\exp(\frac{2\pi i}{n}jk)]_{j,k=1}^{n}$ is a BH$(n,n)$ matrix.
Complex Hadamard matrices have many applications in theoretical quantum physics. Examples include mutually unbiased bases (MUB) [3, 8], quantum teleportation schemes [25] and quantum error correction [9, 16]. They are also used in studies of purely mathematical subjects such as operator algebras [11, 18], equiangular lines [10], and spectral sets [14, 24]. It is also believed that a better understanding of complex Hadamard matrices might shed light on the long-standing Hadamard conjecture of the existence of BH$(n,2)$ matrices.
Complex Hadamard matrices have been classified only for orders less than $6$ [11]. In this article, we present six novel families of complex Hadamard matrices of order 8. Each family contains matrices that are not equivalent to any matrix previously appearing in the literature.
The outline of this paper is as follows. In Section 2, we review some basic theory and definitions, generalise the concept of a palindromic polynomial to multivariate polynomials, and derive a few helpful lemmas. In Section 3, we list complex Hadamard matrices of order 8 already known in the literature. As we will observe, previously presented lists contain imperfections. In Section 4, we present the new families. In Section 5, we establish that the families are pairwise inequivalent by leaning on the results of an exhaustive computer search. Finally, we show that there exist Butson matrices that do not fit in any of the known families.
2 Background material
2.1 Families of complex Hadamard matrices
The general theory and applications of Hadamard matrices are covered in [1] and [12]. The complex case has been studied more comprehensively in [2, 19, 22].
Definition 2.1.
Two complex Hadamard matrices $H_{1}$ and $H_{2}$ are said to be equivalent if there exist permutation matrices $P_{1},P_{2}$ and diagonal matrices $D_{1},D_{2}\in M_{n}(\mathbb{T})$ such that
$H_{1}=D_{1}P_{1}H_{2}P_{2}D_{2}.$
When two matrices are equivalent, we write $H_{1}\cong H_{2}$. If the equality can be written so that $D_{1}=D_{2}=I$, we say that the matrices are permutation equivalent.
A complex Hadamard matrix is called dephased if its first row and first column consist only of ones. We define the dephase operator $\mathscr{D}$ by $\mathscr{D}(H)=D_{1}HD_{2},$ where $D_{1}=\text{diag}(\bar{h}_{1,1},\bar{h}_{2,1},\ldots,\bar{h}_{n,1})$ and $D_{2}=\text{diag}(1,h_{1,1}\bar{h}_{1,2},\ldots,$ $h_{1,1}\bar{h}_{1,n})$. This operator turns an arbitrary matrix into a dephased one, so for every complex Hadamard matrix $H$ there exists an equivalent dephased complex Hadamard matrix $\mathscr{D}(H)$. The dephased form of a matrix is not unique, and there can be equivalent dephased matrices that are not permutation equivalent. The $(n-1)\times(n-1)$ lower right submatrix of a dephased complex Hadamard matrix is called the core. We denote the set of dephased complex Hadamard matrices of order $n$ by $\mathbb{H}_{d}(n)$.
A matrix $H\in\mathbb{H}(n)$ is called isolated if there exists a neighbourhood of $H$ that does not contain matrices that are not equivalent to $H$. We highlight the isolation of matrix $H$ with a superscript notation $H^{(0)}$.
Consider a continuous function $f:D\to\mathbb{H}_{d}(n)$, where $D\subseteq\mathbb{R}^{k}$. Additionally, we require that $k$ coincides with the rank of the Jacobian matrix of $f$ at some point in the interior of $D$. We call the set $H^{(k)}_{n}:=f(D)$ a k-parameter family of complex Hadamard matrices. In practice, it is often more convenient to define the function from $\mathbb{T}^{k}$ to $\mathbb{H}(n)$.
We say that a family of complex Hadamard matrices given by a function $f$ is stemming from a matrix $H\in\mathbb{H}(n)$ if $f(0,\ldots,0)=H$. A family of complex Hadamard matrices is called an affine family if it is an image of a function $f$ in the form $f(x_{1},\ldots,x_{k})=[e^{ig_{j,l}}]_{j,l}$, where $g_{j,l}$ is an affine map on variables $x_{1},\ldots,x_{k}$. A family that is not affine is called a non-affine family.
Let $H^{(k)}_{n}$ and $G^{(k)}_{n}$ be families of complex Hadamard matrices. If for every $A\in H^{(k)}_{n}$ there exists $B\in G^{(k)}_{n}$ such that $A\cong B$ we say that $H^{(k)}_{n}$ is contained in $G^{(k)}_{n}$. If $H^{(k)}_{n}$ is contained in $G^{(k)}_{n}$ and $G^{(k)}_{n}$ is contained in $H^{(k)}_{n}$ we say that the two families are equivalent and write $H^{(k)}_{n}\cong G^{(k)}_{n}$.
For more intuitive notation, we write that $H\in H^{(k)}_{n}$ also in the cases where $H$ is only equivalent to a matrix in $H^{(k)}_{n}$. We define a transpose of a family with $(H_{n}^{(k)})^{T}=\{H^{T}\;|\;H\in H^{(k)}_{n}\}$. We call a family $H^{(k)}$ symmetric if $H_{n}^{(k)}\cong(H_{n}^{(k)})^{T}.$
For each $H\in\mathbb{H}(n)$ a defect-value [23], notated by $d(H)$, can be computed. A defect value of a matrix gives an upper bound on the dimension of the smooth manifold in $\mathbb{H}(n)$ to which it belongs. In particular, $d(H)=0$ implies that $H$ is isolated.
2.2 Palindromic polynomials
Reciprocal and palindromic polynomials are usually defined as univariate polynomials, but we generalise the definition for multivariate polynomials as well.
Definition 2.2.
A conjugate reciprocal polynomial $f^{*}$ of a multivariate polynomial $f\in\mathbb{C}[x_{1},\ldots,x_{n}]$ is a polynomial
$f^{*}(x_{1},\ldots,x_{n})=\overline{f(\overline{x_{1}}^{-1},\ldots,\overline{x_{n}}^{-1})}.$
If the coefficients of the polynomial $f(x_{1},\ldots,x_{n})$ are real, then Definition 2.2 coincides with the usual definition of reciprocal polynomial $f^{*}(x_{1},\ldots,x_{n})=f(1/x_{1},\ldots,1/x_{n})$. If the variables are taken to be unimodular, then it coincides with the conjugate $f^{*}(x_{1},\ldots,x_{n})=\overline{f(x_{1},\ldots,x_{n})}$.
Definition 2.3.
A multivariate polynomial $f\in\mathbb{C}[x_{1},\ldots,x_{n}]$ is called palindromic if
$f=x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}f^{*}$
and anti-palindromic if
$f=-x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}f^{*},$
where $\alpha_{i}$ is the maximum degree of $x_{i}$ in $f(x_{1},\ldots,x_{n})$.
The symmetry of palindromic and anti-palindromic polynomials can often simplify the process of solving systems involving them.
Lemma 2.1.
Let $p=\sum_{j=1}^{l}c_{j}\mathbf{x_{j}}\in\mathbb{C}[x_{1},\ldots,x_{k}]$ be palindromic or anti-palindromic with an even number of monomials $\mathbf{x_{j}}$ in lexicographic order. Then the roots of $p$ with all variables nonzero correspond to solutions of
$\sum_{j=0}^{l/2}(c_{j}y_{j}\pm\overline{c_{j}}y_{j}^{-1})=0,$
where a positive sign corresponds to the palindromic case, a negative sign to the anti-palindromic case and $y_{j}=\mathbf{x_{j}}/(x_{1}^{\alpha_{1}/2}\cdots x_{k}^{\alpha_{k}/2})$.
Proof.
Comparing the terms in
$(x_{1}^{-\alpha_{1}/2}\cdots x_{k}^{-\alpha_{k}/2})\sum_{j=0}^{l}c_{j}\mathbf{x_{j}}=(x_{1}^{\alpha_{1}/2}\cdots x_{k}^{\alpha_{k}/2})\sum_{j=0}^{l}\overline{c_{j}}\mathbf{x_{j}}^{-1}$
gives the result. ∎
Lemma 2.1 also provides a method for expressing unimodular roots of palindromic polynomials with unimodular coefficients as real roots of trigonometric functions. After the change of variables $x_{j}\to e^{i\xi_{j}}$, we can employ the identities $2\cos(\xi)=e^{i\xi}+e^{-i\xi}$ and $2i\sin(\xi)=e^{i\xi}-e^{-i\xi}$. For example, finding solutions numerically can be more convenient with real numbers.
We will be interested in the unimodular roots of palindromic polynomials, and the following lemmas provide some bounds on when the obtained solution is unimodular.
Lemma 2.2.
Let $f$ be a multivariate palindromic polynomial of $n$ variables, where $\alpha_{1}=1$, that is,
$f(x_{1},\ldots,x_{n})=p_{1}(x_{2},\ldots,x_{n})x_{1}+p_{0}(x_{2},\ldots,x_{n}).$
Fix $x_{2},\ldots,x_{n}\in\mathbb{T}$. If $f(x_{1},\ldots,x_{n})=0$ and $p_{1}(x_{2},\ldots,x_{n})\neq 0$ then $x_{1}\in\mathbb{T}$.
Proof.
The palindromic property gives $p_{1}=\overline{p_{0}}\cdot x_{2}^{\alpha_{2}}\cdots x^{\alpha_{n}}_{n}$. Hence, solving $x_{1}$ gives $x_{1}=-p_{0}/p_{1}=-p_{0}/\overline{p_{0}}\cdot x_{2}^{-\alpha_{2}}\cdots x^{-\alpha_{n}}_{n}$ which has an absolute value of 1. ∎
Lemma 2.3.
Let $f$ be a multivariate palindromic polynomial of $n$ variables, where $\alpha_{1}=2$, that is,
$f(x_{1},\ldots,x_{n})=p_{2}(x_{2},\ldots,x_{n})x^{2}_{1}+p_{1}(x_{2},\ldots,x_{n})x_{1}+p_{0}(x_{2},\ldots,x_{n}).$
Fix $x_{2},\ldots,x_{n}\in\mathbb{T}$. If $f(x_{1},\ldots,x_{n})=0$ and $p_{2}(x_{2},\ldots,x_{n})\neq 0$ then $x_{1}\in\mathbb{T}$ if and only if $|p_{1}|/|p_{2}|\leq 2$.
Proof.
The palindromic property gives $p_{2}=\overline{p_{0}}\cdot x_{2}^{\alpha_{2}}\cdots x^{\alpha_{n}}_{n}$ and $p_{1}=\overline{p_{1}}\cdot x_{2}^{\alpha_{2}}\cdots x^{\alpha_{n}}_{n}.$ Notice that $x_{2}^{\alpha_{2}}\cdots x^{\alpha_{n}}_{n}=e^{2\arg(p_{1})i}.$ We can solve $x$ from
$|p_{2}|e^{\arg(p_{2})i}x_{1}^{2}+|p_{1}|e^{\arg(p_{1})i}x_{1}+|p_{2}|e^{(2\arg(p_{1})-\arg(p_{2}))i}=0.$ (2.1)
Division with the constant term and a variable change $x_{1}=ye^{(\arg(p_{2})-\arg(p_{1}))}$ gives
$y^{2}+\frac{|p_{1}|}{|p_{2}|}y+1=0.$ (2.2)
Unimodular solutions of $y$ are exactly the unimodular solutions of $x$. We get that
$y=\frac{1}{2}\left(-|p_{1}|/|p_{2}|\pm i\sqrt{4-|p_{1}|^{2}/|p_{2}|^{2}}\right),$
which is unimodular if and only if $|p_{1}|/|p_{2}|\leq 2$. ∎
Notice that the discriminant of (2.1) is the discriminant of (2.2) scaled with $x_{2}^{\alpha_{2}}\cdots x^{\alpha_{n}}_{n}|p_{2}|$. In practice it can be more convenient to consider the inequality $C/(x_{2}^{\alpha_{2}}\cdots x^{\alpha_{n}}_{n})\leq 0$ instead of $|p_{1}|/|p_{2}|\leq 2$, where $C$ is the discriminant of (2.1). Moreover, Lemma 2.2 and Lemma 2.3 can be shown to work for anti-palindromic polynomials as well with almost identical proofs.
3 Known complex Hadamard matrices of order 8
3.1 Affine families
In the previous literature, the following affine families of order 8 have appeared
$F^{(5)}_{8},D^{(5)}_{8A},\;D^{(5)}_{8B},\;S_{8A}^{(4)},\;S^{(4)}_{8B}.$ (3.1)
The family $F_{8}^{(5)}:=F_{8}(\mathbb{T}^{5})$, introduced in [22], is 5-parameter family stemming from the Fourier matrix where $F_{8}(a,b,c,d,e)=$
$\begin{bmatrix}1&1&1&1&1&1&1&1\\ 1&\omega&\omega^{2}&\omega^{3}&\omega^{4}&\omega^{5}&\omega^{6}&\omega^{7}\\ 1&\omega^{2}&\omega^{4}&\omega^{6}&1&\omega^{2}&\omega^{4}&\omega^{6}\\ 1&\omega^{3}&\omega^{6}&w&\omega^{4}&\omega^{7}&\omega^{2}&\omega^{5}\\ 1&\omega^{4}&1&\omega^{4}&1&\omega^{4}&1&\omega^{4}\\ 1&\omega^{5}&\omega^{2}&\omega^{7}&\omega^{4}&w&\omega^{6}&\omega^{3}\\ 1&\omega^{6}&\omega^{4}&\omega^{2}&1&\omega^{6}&\omega^{4}&\omega^{2}\\ 1&\omega^{7}&\omega^{6}&\omega^{5}&\omega^{4}&\omega^{3}&\omega^{2}&\omega\end{bmatrix}\circ\begin{bmatrix}1&1&1&1&1&1&1&1\\ 1&a&b&c&1&a&b&c\\ 1&d&1&d&1&d&1&d\\ 1&e&b&\overline{a}ce&1&e&b&\overline{a}ce\\ 1&1&1&1&1&1&1&1\\ 1&a&b&c&1&a&b&c\\ 1&d&1&d&1&d&1&d\\ 1&e&b&\overline{a}ce&1&e&b&\overline{a}ce\end{bmatrix},$
$\omega=\exp(2\pi i/8)$ and $\circ$ is the Hadamard product.
The family $D_{8A}^{(5)}:=D_{8A}(\mathbb{T}^{5})$, introduced in [7], where $D_{8A}(a,b,c,d,e)=$
$\begin{bmatrix}1&1&1&1&1&1&1&1\\ 1&\omega&\omega^{2}&\omega^{3}&\omega^{4}&\omega^{5}&\omega^{6}&\omega^{7}\\ 1&\omega^{2}&\omega^{6}&\omega^{4}&1&\omega^{2}&\omega^{6}&\omega^{4}\\ 1&\omega^{3}&\omega^{6}&\omega&\omega^{4}&\omega^{7}&\omega^{2}&\omega^{5}\\ 1&\omega^{4}&\omega^{4}&1&1&\omega^{4}&\omega^{4}&1\\ 1&\omega^{5}&\omega^{2}&\omega^{7}&\omega^{4}&\omega&\omega^{6}&\omega^{3}\\ 1&\omega^{6}&\omega^{2}&\omega^{4}&1&\omega^{6}&\omega^{2}&\omega^{4}\\ 1&\omega^{7}&\omega^{6}&\omega^{5}&\omega^{4}&\omega^{3}&\omega^{2}&\omega,\end{bmatrix}\circ\begin{bmatrix}1&1&1&1&1&1&1&1\\ 1&a&b&c&1&a&b&c\\ 1&d&1&d&1&d&1&d\\ 1&e&b&\overline{a}ce&1&e&b&\overline{a}ce\\ 1&1&1&1&1&1&1&1\\ 1&a&b&c&1&a&b&c\\ 1&d&1&d&1&d&1&d\\ 1&e&b&\overline{a}ce&1&e&b&\overline{a}ce\end{bmatrix}$
has a very similar structure to $F^{(5)}_{8}$, but unlike $F^{(5)}_{8}$, it is not symmetric. The family $D_{8B}^{(5)}$ is the transpose of the family $D_{8A}^{(5)}$.
Finally, the 4-parameter family $S_{8A}^{(4)}:=S_{8A}(\mathbb{T}^{4})$, introduced in [15], is given by $S_{8A}(a,b,c,d)=$
$\begin{bmatrix}1&1&1&1&1&1&1&1\\ 1&1&-1&-1&-1&i&-i&1\\ 1&i&i&-i&1&-1&-1&-i\\ 1&i&-i&i&-1&-i&i&-i\\ 1&-1&-i&i&1&i&-i&-1\\ 1&-1&i&-i&-1&1&1&-1\\ 1&-i&-1&-1&1&-i&i&i\\ 1&-i&1&1&-1&-1&-1&i\\ \end{bmatrix}\circ\begin{bmatrix}1&1&1&1&1&1&1&1\\ 1&d&d&d&1&cd&cd&d\\ 1&a\overline{d}&b\overline{d}&b\overline{d}&1&1&1&a\overline{d}\\ 1&a&b&b&1&cd&cd&a\\ 1&1&b\overline{d}&b\overline{d}&1&c&c&1\\ 1&d&b&b&1&d&d&d\\ 1&a\overline{d}&1&1&1&c&c&a\overline{d}\\ 1&a&d&d&1&d&d&a\end{bmatrix}$
and its transpose is denoted by $S_{8B}^{(4)}\cong(S_{8A}^{(4)})^{T}.$ The family $S_{8A}^{(4)}$ is often referred to without the subscript $A$ and its transpose is sometimes called $D^{(4)}_{8}$.
As far as we know, the pairwise relationships among these families have not all been explored. At least families $D_{8A}^{(5)}$ and $D_{8B}^{(5)}$ are missing from some of the earlier listings of affine families of order 8.
3.2 Isolated matrices
Four isolated complex Hadamard matrices of order 8 are known, namely
$A_{8A}^{(0)},\;A^{(0)}_{8B},\;V_{8A}^{(0)},\;V_{8B}^{(0)}.$ (3.2)
The matrix
$A_{8A}^{(0)}=\begin{bmatrix}a&1&-a&1&-a&-a&a&1\\ 1&-a&1&a&-a&a&-a&1\\ -a&1&-a&a&1&a&1&-a\\ 1&a&a&1&1&-a&-a&-a\\ -a&-a&1&1&a&-a&1&a\\ -a&a&a&-a&-a&1&1&1\\ a&-a&1&-a&1&1&a&-a\\ 1&1&-a&-a&a&1&-a&a\\ \end{bmatrix},$
where $a$ is the root of the polynomial $33x^{2}-2x+$ 3, is introduced in [4]. The matrix $A^{(0)}_{8A}$ is symmetric but not equivalent to its conjugate, which is denoted by $A_{8B}^{(0)}$.
Consider a matrix-valued function
$V_{8}(a,b,c)=\begin{bmatrix}-ab&-ab&bc&bc&1&1&ac&ac\\ -ab&bc&-ab&1&bc&ac&1&-ac\\ bc&-ab&1&-ab&ac&bc&-ac&1\\ bc&1&-ab&ac&-ab&-ac&bc&-1\\ 1&bc&ac&-ab&-ac&-ab&-1&bc\\ 1&ac&bc&-ac&-ab&-1&-ab&-bc\\ ac&1&-ac&bc&-1&-ab&-bc&-ab\\ ac&-ac&1&-1&bc&-bc&-ab&ab\\ \end{bmatrix}.$
The matrices $V_{8A}^{(0)}$ and $V_{8B}^{(0)}$ are isolated matrices, presented in [6] and analytically solved later in [5]. These matrices are found by solving the system of polynomial equations arising from orthogonality constrains of the matrix $V_{8}(\frac{\sqrt{a}}{\sqrt{b}\sqrt{c}},\frac{\sqrt{b}}{\sqrt{a}\sqrt{c}},\frac{\sqrt{a}\sqrt{b}}{\sqrt{c}})$. In [5], analytic solutions for these equations are obtained from roots of quartic polynomials. We present an alternative solution which leads to a more compact way of expressing the solutions with trigonometric functions. In some previous literature, all the solutions have been treated as different matrices, even though there are only two that are inequivalent.
Lemma 3.1.
$V_{8}(a,b,c)$ is a complex Hadamard matrix if and only if
[…]
where $a,b,c\in\mathbb{T}$.
Proof.
Orthogonality requirements between the rows of $V_{8}(a,b,c)$ give us a system of polynomial equations $p_{V}^{(1)}=0$, $p_{V}^{(2)}=0$ and $p^{(3)}_{V}=0$, where
$p_{V}^{(1)}=b+a^{2}b+c-a^{2}c-b^{2}c+a^{2}b^{2}c+bc^{2}+a^{2}bc^{2},$
$p_{V}^{(2)}=a+b-a^{2}b-bc^{2}+a^{2}bc^{2}+ab^{2}c^{2},$
$p^{(3)}_{V}=ab^{2}+c+a^{2}c+b^{2}c+a^{2}b^{2}c+ac^{2}.$
We can equivalently consider a system $p^{(1)}_{V}=0$, $p^{(2)}_{V}+p_{V}^{(3)}=0$ and $p^{(2)}_{V}-p_{V}^{(3)}=0$. Notice that $p^{(1)}_{V}$, $p^{(2)}_{V}+p_{V}^{(3)}$ and $p^{(2)}_{V}-p_{V}^{(3)}$ are palindromic. Consider the constraint $p^{(2)}_{V}+p_{V}^{(3)}=0$. Lemma 2.1 gives an equivalent constraint
$ab+\frac{1}{ab}+\frac{a}{b}+\frac{b}{a}+ac+\frac{1}{ac}-\frac{a}{c}-\frac{c}{a}+bc+\frac{1}{bc}+\frac{c}{b}+\frac{b}{c}=0,$
which factors to
$\left(a+\frac{1}{a}\right)\left(b+\frac{1}{b}\right)+\left(a-\frac{1}{a}\right)\left(c-\frac{1}{c}\right)+\left(b+\frac{1}{b}\right)\left(c+\frac{1}{c}\right)=0$
and by unimodularity of $a,b,c$ we have
[…]
Other constraints can be dealt with in a similar manner. ∎
Proposition 3.1.
Let
$x_{1}=\sqrt{\frac{1}{2}\sin\left(\frac{\pi}{16}\right)\sec\left(\frac{3\pi}{16}\right)},\quad$ $x_{2}=\sqrt{\cos\left(\frac{\pi}{16}\right)\cos\left(\frac{3\pi}{16}\right)\sec\left(\frac{\pi}{8}\right)},$
$y_{1}=2\cos\left(\frac{\pi}{16}\right)\sqrt{\sin\left(\frac{\pi}{16}\right)\sec\left(\frac{3\pi}{16}\right)},$
$y_{2}=-\sin\left(\frac{\pi}{16}\right)\sqrt{2\cos\left(\frac{\pi}{16}\right)\sec\left(\frac{\pi}{8}\right)\sec\left(\frac{3\pi}{16}\right)},$
$z_{1}=-2\cos\left(\frac{\pi}{16}\right)\sec\left(\frac{\pi}{8}\right)\sqrt{\sin\left(\frac{\pi}{16}\right)\cos\left(\frac{3\pi}{16}\right)},$
$z_{2}=-\frac{1}{2}\sec\left(\frac{\pi}{8}\right)\sqrt{\cos\left(\frac{3\pi}{16}\right)\sec\left(\frac{\pi}{16}\right)\sec\left(\frac{\pi}{8}\right)}.$
$V_{8}(x_{1}+x_{2}i,\;y_{1}+y_{2}i,\;z_{1}+z_{2}i)$ is a complex Hadamard matrix.
Proof.
The given assignment is a solution for the system of Lemma 3.1. ∎
We define $V_{8B}^{(0)}:=V_{8}(x_{1}+x_{2}i,\;y_{1}+y_{2}i,\;z_{1}+z_{2}i)$ and $V_{8A}^{(0)}:=(V_{8B}^{(0)})^{\dagger}$. Let $P_{1}=(e_{4}\;e_{3}\;e_{7}\;e_{8}\;e_{1}\;e_{2}\;e_{6}\;e_{5})$ and $P_{2}=(e_{1}\;e_{7}\;e_{6}\;e_{4}\;e_{5}\;e_{3}\;e_{2}\;e_{8})$. The matrices $\mathscr{D}(V_{8A}^{(0)})$ and $\mathscr{D}(P_{1}V_{8B}^{(0)}P_{2})$ matches exactly the numerical matrices of the same name given in [6].
All the 16 solutions to the system of Lemma 3.1 that do not produce Butson matrices have the form
$(a,b,c)=(\mu_{1}\cdot x_{1}+\mu_{2}\cdot x_{2}i,\;\mu_{3}\cdot y_{1}+\mu_{2}\cdot y_{2}i,\;\mu_{1}\cdot z_{1}+\mu_{1}\mu_{2}\mu_{3}\cdot z_{2}i)$
or
$(a,b,c)=(\mu_{1}\cdot x_{2}+\mu_{2}\cdot x_{1}i,\;\mu_{3}\cdot z_{2}+\mu_{2}\cdot z_{1}i,\;\mu_{1}\cdot y_{2}+\mu_{1}\mu_{2}\mu_{3}\cdot y_{1}i),$
where $\mu_{1},\mu_{2},\mu_{3}\in\{-1,1\}$. These can all be shown to produce matrices that are equivalent to either $V_{8B}^{(0)}$ or $V_{8A}^{(0)}$. The numerical values for the different solutions given in [6] are a perfect match to the approximate values of $(\frac{c}{b},\frac{c}{a},\frac{1}{ab})$, where $(a,b,c)$ is one of the solutions above.
3.3 Non-affine families
The existing literature also knows two non-affine families:
$T^{(1)}_{8},\;T_{8B}^{(3)}.$ (3.3)
The first is a 1-parameter family introduced in [4]. The family is given by function $T_{8}(x,y,z,u)=$
$\begin{bmatrix}1&1&1&1&1&1&1&1\\ 1&1&1&1&1&\zeta^{3}&\zeta^{15}&\zeta^{18}\\ 1&\zeta^{8}&\zeta^{8}&\zeta^{8}&\zeta^{8}&\zeta^{13}&\zeta^{15}&\zeta^{8}\\ 1&\zeta^{2}&\zeta^{2}&\zeta^{2}&\zeta^{2}&\zeta^{17}&\zeta^{5}&\zeta^{2}\\ 1&1&1&1&1&\zeta^{7}&\zeta^{5}&\zeta^{12}\\ 1&1&1&1&1&1&\zeta^{10}&\zeta^{10}\\ 1&1&1&1&1&\zeta^{10}&1&\zeta^{10}\\ 1&1&1&1&1&\zeta^{10}&\zeta^{10}&1\end{bmatrix}\circ\begin{bmatrix}1&1&1&1&1&1&1&1\\ 1&x&y&z&-\frac{yz}{x}&1&1&1\\ 1&-\frac{u}{y}&\frac{u}{x}&-\frac{uz}{xy}&-\frac{uz}{x^{2}}&1&1&1\\ 1&y&x&-\frac{yz}{x}&z&1&1&1\\ 1&\frac{u}{x}&-\frac{u}{y}&-\frac{uz}{x^{2}}&-\frac{uz}{xy}&1&1&1\\ 1&-i\frac{x}{z}&i\frac{x}{z}&-i\frac{z}{x}&i\frac{z}{x}&1&1&1\\ 1&u&-u&-\frac{uz^{2}}{x^{2}}&\frac{uz^{2}}{x^{2}}&1&1&1\\ 1&i\frac{uz}{x}&i\frac{uz}{x}&-i\frac{uz}{x}&-i\frac{uz}{x}&1&1&1\\ \end{bmatrix},$
where $\zeta=e^{2\pi i/20}$. The domain where the function $T(x,y,z,u)$ gives Hadamard matrices can be expressed solely in terms of $x$. The relations of the other variables and $x$ are complicated and thus not presented here. The parameter $x$ is not free on $\mathbb{T}$, and the exact bounds that guarantee the unimodularity of the other variables are not known. We will show in Section 4 that this is most likely a subfamily of a new family we have discovered.
The non-affine 3-parameter family $T^{(3)}_{8B}$ is introduced in [5] and given by the function $T_{8B}(a,b,c,d)=$
$\begin{bmatrix}1&1&1&1&1&1&1&1\\ 1&-1&-d&-c&d&c&cd&-cd\\ 1&-a&-1&c&a&-ac&-c&ac\\ 1&a&d&-1&ad&-a&-d&-ad\\ 1&b&-b&-bc&-1&-c&c&bc\\ 1&-b&bd&b&-d&-1&d&-bd\\ 1&-ab&b&-b&-a&a&-1&ab\\ 1&ab&-bd&bc&-ad&ac&-cd&-abcd\\ \end{bmatrix},$
with some additional constraints for $d$. As the family is discussed only briefly in [5], we formulate the following more detailed result.
Proposition 3.2.
Let $T_{8B}^{(3)}:=\mathcal{S}_{B}(\mathbb{T}^{3}\backslash D_{0})$, where
$\mathcal{S}_{B}(a,b,c)=T_{8B}(a,b,c,d^{\prime})$
with
$d^{\prime}=\frac{1+ab+ac+bc}{a+b+c+abc}$
and $D_{0}=\{(a,b,c)\in\mathbb{T}^{3}\;|\;-1,1\in\{a,b,c\}\}$. $T_{8B}^{(3)}$ is a symmetric 3-parameter family of complex Hadamard matrices. Additionally, for every $H\in T_{8B}(\mathbb{T}^{4})\cap\mathbb{H}(8)$ there exists an equivalent matrix in $T^{(3)}_{8B}.$
Proof.
Each orthogonality constraint between the rows of $T_{8B}$ simplifies to 0 or
$p_{B}=1+ab+ac+bc-d(a+b+c+abc)=0,$ (3.4)
from which $d$ can be solved with respect to three other variables. We can notice that $p_{B}$ is palindromic. By Lemma 2.2, the obtained solution for $d$ gives unimodular values for any $a,b,c\in\mathbb{T}$ as long as the denominator is not 0. The denominator is 0 exactly on the set $D_{0}$.
Consider permutation matrices $P_{1}=(e_{1}\;e_{2}\;e_{4}\;e_{6}\;e_{3}\;e_{5}\;e_{7}\;e_{8})$ and $P_{2}=(e_{1}\;e_{2}\;e_{5}\;e_{3}\;e_{6}\;e_{4}\;e_{7}\;e_{8})$. We have that $\mathcal{S}_{B}(-d^{\prime},c,a)=P_{1}\mathcal{S}_{B}^{T}P_{2}(a,b,c)$ implying that $T^{(3)}_{8B}$ is symmetric. Additionally, one can find similar assignments of $\mathcal{S}_{B}$ that show that the Hadamard matrices in $T_{8B}(D_{0})$ are all equivalent to some matrix in $T^{(3)}_{8B}$. ∎
In [5], a third non-affine family called $T^{(3)}_{8C}$ is also introduced. We have left this family out of our considerations as the relationships of the entries are not fully solved, and its existence is observed by utilising numerical methods. We have, however, found and solved a family that seems to contain all the matrices that are obtained from the given numerical construction of $T^{(3)}_{8C}$. We will be referring to it with the same name.
4 Novel matrix families
In this section, we introduce four new families of complex Hadamard matrices. All the families are given as the image of a matrix-valued function on a specific domain.
To be precise, we should show that there are points in the domains to be presented. We have left that to Section 5, where we show that there are points in the domains that produce matrices that do not belong to any of the other families.
[…]
4.1 Family $T_{8C}^{(3)}$
Consider a matrix-valued function $T_{8C}(a,b,c,d,e,f)=$
$\begin{bmatrix}a&bd&d&ab&ad&bc&ac&bd\\ bd&-a&ab&-d&bc&-ad&bd&-ac\\ d&ab&bdef&-aef&acf&bdf&bce&-ade\\ ab&-d&-aef&-bdef&bdf&-acf&-ade&-bce\\ ac&bd&adf&bcf&-cdf&-abcf&-ac&-bcd\\ bd&-ac&bcf&-adf&-abcf&cdf&-bcd&ac\\ ad&bc&bde&-ace&-ac&-bcd&-abce&cde\\ bc&-ad&-ace&-bde&-bcd&ac&cde&abce\\ \end{bmatrix}.$
Lemma 4.1.
$T_{8C}(a,b,c,d,e,f)$ is a complex Hadamard matrix if and only if $p^{(1)}_{C}=0$, $p^{(2)}_{C}=0$ and $p^{(3)}_{C}=0$, where
$p^{(1)}_{C}=$ $a^{2}c-a^{2}b^{2}c+a^{2}c^{2}-b^{2}c^{2}+a^{2}d^{2}-b^{2}d^{2}+cd^{2}-b^{2}cd^{2},$
$p^{(2)}_{C}=$ $a^{2}bc+bcd^{2}-a^{2}cde+b^{2}cde+abc^{2}f+abd^{2}f-acdef+ab^{2}cdef,$
$p^{(3)}_{C}=$ $acd-ab^{2}cd+abc^{2}e+abd^{2}e+a^{2}cdf-b^{2}cdf+a^{2}bcef+bcd^{2}ef$
and $a,b,c,d,e,f\in\mathbb{T}.$
Proof.
We need that $a,b,c,d,e,f\in\mathbb{T}$ for unimodular entries. Now we can identify complex conjugates of the variables with their multiplicative inverses. Each inner product between two distinct rows of $T_{8C}$, that is not equal to 0, can be factored to $p_{C}^{(i)}/q$ for some $i\in\{1,2,3\}$, where $q$ is a monomial on variables $a,b,c,d,e,f$. ∎
Theorem 4.1.
Let $T^{(3)}_{8C}:=S_{C}(D)$, where
$\mathcal{S}_{C}(a,b,c)=T_{8C}(a,b,c,d^{\prime},e^{\prime},f^{\prime}),$
with
$d^{\prime}=$ $-\sqrt{cA_{1}}/\sqrt{B_{1}},$
$e^{\prime}=$ $d^{\prime}((1-a)(1+a)(a-b^{2})(a+b^{2})A_{3}+\sqrt{C})/B_{3},$
$f^{\prime}=$ $((c-b^{2})(c+1)(a^{2}+c)(a^{2}-b^{2}c)A_{2}+\sqrt{C})/B_{2},$
where
$A_{1}=$ $-a^{2}+a^{2}b^{2}-a^{2}c+b^{2}c,$
$A_{2}=$ $a^{2}+a^{4}-4a^{2}b^{2}+b^{4}+a^{2}b^{4},$
$A_{3}=$ $-a^{4}b^{2}-a^{4}c+a^{2}b^{2}c+a^{4}b^{2}c-a^{2}b^{4}c-a^{2}c^{2}-a^{4}c^{2}+6a^{2}b^{2}c^{2}$
$-b^{4}c^{2}-a^{2}b^{4}c^{2}-a^{2}c^{3}+b^{2}c^{3}+a^{2}b^{2}c^{3}-b^{4}c^{3}-b^{2}c^{4},$
$B_{1}=$ $a^{2}-b^{2}+c-b^{2}c,$
$B_{2}=$ $-2a(a-b)(-1+b)(1+b)(a+b)A_{3},$
$B_{3}=$ $2bc(a-c)(a+c)A_{1}A_{2},$
$C=$ $(1+c)^{2}(a^{2}+c)^{2}(-b^{2}+c)^{2}(a^{2}-b^{2}c)^{2}A_{2}^{2}-4a^{2}(a^{2}-b^{2})^{2}(-1+b^{2})^{2}A_{3}^{2}$
and
$D=\{(a,b,c)\in\mathbb{T}^{3}\;|\;C/(a^{8}b^{8}c^{4})\leq 0,\;B_{1},B_{2},B_{3}\neq 0\}.$
$T^{(3)}_{8C}$ is a 3-parameter family of complex Hadamard matrices.
Proof.
One can check that the assignment $d=d^{\prime}$, $e=e^{\prime}$, and $f=f^{\prime}$ satisfy the system in Lemma 4.1. We can notice that $e^{\prime}$ and $f^{\prime}$ are roots of a palindromic polynomial, so by Lemma 2.3 they get unimodular values whenever $a,b,c,d^{\prime}\in\mathbb{T}$ and $C/(a^{8}b^{8}c^{4})\leq 0$. Also, $d^{\prime}$ is a root of an anti-palindromic polynomial, and it gets unimodular values whenever $a,b,c\in\mathbb{T}$. ∎
There are other solutions to the system in Lemma 4.1 than the one given in Theorem 4.1. The variables $(d,e,f)$ could also be assigned to $(-d^{\prime},-e^{\prime},f^{\prime})$, $(d^{\prime},-\overline{e^{\prime}},\overline{f^{\prime}})$ or $(-d^{\prime},\overline{e^{\prime}},\overline{f})$ corresponding different branches of the square root. Numerical results suggest that the different solutions result in equivalent families. There are also complex Hadamard matrices that correspond to roots of $B_{1},$ $B_{2}$ or $B_{3}$ and thus are not included in the family. One such matrix is the unique real Hadamard matrix of order 8.
It can be noticed that if $(a,b,c,d,e,f)$ is a solution to the system in Lemma 4.1, then so is $(a,b,d^{2}/c,d,e,f)$. We also have that $\mathscr{D}(T_{8C})^{T}(a,b,c,d,e,f)=\mathscr{D}(T_{8C})(a,b,d^{2}/c,d,e)$. However, it does not always hold that $\mathscr{D}(\mathcal{S}_{C})^{T}(a,b,c)=\mathscr{D}(\mathcal{S}_{C})(a,b,(d^{\prime})^{2}/c)$, which would imply that $T_{8C}^{(3)}$ is symmetric, as the solution $(a,b,d^{2}/c,d,e,f)$ can also be in one of the other solution branches. On the other hand, as the different branches seem to give equivalent families, we conjecture that $T^{(3)}_{8C}$ is symmetric.
4.2 Family $T_{8D}^{(3)}$
Consider a matrix-valued function $T_{8D}(a,b,c,d,e,f)=$
$\begin{bmatrix}1&-c&c&-e&-c&e&-e&ce\\ 1&c&-e&c&-e&c&-e&-ce\\ c&-1&c&e&-c&-e&ce&-e\\ c&1&e&c&e&c&ce&e\\ b&f&ad&ab&df&bf&abd&adf\\ b&-f&ab&ad&-bf&-df&abd&-adf\\ f&b&-ad&ab&-df&bf&-adf&-abd\\ f&-b&ab&-ad&-bf&df&-adf&abd\\ \end{bmatrix}.$
Lemma 4.2.
$T_{8D}(a,b,c,d,e,f)$ is a complex Hadamard matrix if and only if $p^{(1)}_{D}=0$, $p^{(2)}_{D}=0$, $p^{(3)}_{D}=0$ and $p^{(4)}_{D}=0$, where
$p^{(1)}_{D}=$ $abc+abcd-bce-ade-bcf-adf+ef+def,$
$p^{(2)}_{D}=$ $\;acd+abcd-abe-bce+adf+cdf-ef-bef,$
$p^{(3)}_{D}=$ $\;abc+abd+be+ade-bcf-acdf-cef-def,$
$p^{(4)}_{D}=$ $\;abd+acd+be+abe+cdf+acdf+bef+cef,$
and $a,b,c,d,e,f\in\mathbb{T}.$
Proof.
We need that $a,b,c,d,e,f\in\mathbb{T}$ for unimodular entries. Similarly to Lemma 4.1, the system of equations that requires orthogonality of the rows can be reduced to $p^{(1)}_{D}=0$, $p^{(2)}_{D}=0$, $p^{(3)}_{D}=0$ and $p^{(4)}_{D}=0$. ∎
Theorem 4.2.
Let $T^{(3)}_{8D}:=\mathcal{S}_{D}(D),$ where
$\mathcal{S}_{D}(a,b,c)=T_{8D}(a,b,c,d^{\prime},e^{\prime},f^{\prime})$
with
$d^{\prime}=-A_{1}/B_{1},$
$e^{\prime}=c((a-b)(ab-c)(1+c)+\sqrt{C})/B_{2},$
$f^{\prime}=-((1+c)A_{2}+\sqrt{C})/(2B_{3}),$
where
$A_{1}$ $=bc+abc+b^{2}c+bc^{2},$
$A_{2}$ $=2ab+a^{2}b+ab^{2}+ac+bc+2abc,$
$B_{1}$ $=ab+ac+bc+abc,$
$B_{2}$ $=2(c-1)B_{1},$
$B_{3}$ $=ab+c+2ac+2bc+abc+c^{2},$
$C$ $=(1+c)^{2}A_{2}^{2}-4abB_{3}^{2},$
and
$D=\{(d,e,f)\in\mathbb{T}^{3}\;|\;C/(a^{2}b^{2}c^{2})\leq 0,\;B_{1},B_{2},B_{3}\neq 0\}.$
$T^{(3)}_{8D}$ is a 3-parameter family of complex Hadamard matrices.
Proof.
One can check that the assignment $d=d^{\prime}$, $e=e^{\prime}$, and $f=f^{\prime}$ satisfy the system in Lemma 4.2. Additionally, $d^{\prime}$, $e^{\prime}$ and $f^{\prime}$ are all roots of palindromic polynomials. By Lemma 2.2 and Lemma 2.3, $d^{\prime}$, $e^{\prime}$ and $f^{\prime}$ are unimodular whenever $a,b,c\in\mathbb{T}$ and $C/(a^{2}b^{2}c^{2})\leq 0$. ∎
Changing the sign of the square root also gives a valid solution. Numerical results suggest that this solution results in an equivalent family. There are also multiple solutions for the system in Lemma 4.2 that correspond to points that are roots of $B_{1}$, $B_{2}$ or $B_{3}$ and therefore not included in the solutions given.
In [5], it was suggested that the family $T^{(1)}_{8}$ could be contained in $T_{8B}^{(3)}$ or $T_{8C}^{(3)}$. Testing this with the algorithm given in Section 5.1 shows that this is not the case. The following propositions show that $T^{(1)}_{8}$ is most likely contained in $(T^{(3)}_{8D})^{T}$.
Proposition 4.1.
For all $H\in T^{(1)}_{8}$ there exists $H^{\prime}\in T_{8D}(\mathbb{T}^{6})$ such that $H^{T}\cong H^{\prime}$.
Proof.
Consider permutation matrices
$P_{1}=(e_{1}\;e_{6}\;e_{5}\;e_{8}\;e_{7}\;e_{4}\;e_{3}\;e_{2})\text{ and }P_{2}=(e_{1}\;e_{6}\;e_{3}\;e_{5}\;e_{2}\;e_{4}\;e_{8}\;e_{7}).$
We have that
$\mathscr{D}(T_{8D})(e^{i\pi\frac{3}{10}}u/y,\;iz,\;i,\;-iyz/x,\;e^{i\pi\frac{3}{10}},\;i,ix)=P_{1}T_{8}^{T}P_{2}(x,y,z,y).$
4.3 Family $T_{8E}^{(3)}$
Consider a matrix valued function $T_{8E}(a,b,c,d,e,f)=$
$\begin{bmatrix}1&1&1&1&1&1&1&1\\ 1&d&e&-f&df&ef&-de&def\\ 1&a&-b&f&-af&bf&ab&abf\\ 1&-d&b&c&cd&-bc&bd&bcd\\ 1&-a&-e&-c&-ac&-ce&-ae&ace\\ 1&-ad&be&-1&ad&-be&-abde&abde\\ 1&ad&-1&-cf&-acdf&cf&-ad&acdf\\ 1&-1&-be&cf&-cf&-bcef&be&bcef\\ \end{bmatrix}.$
Lemma 4.3.
$T_{8E}(a,b,c,d,e,f)$ is a complex Hadamard matrix if and only if $p^{(1)}_{E}=0$, $p^{(2)}_{E}=0$, $p^{(3)}_{E}=0$ and $p^{(4)}_{E}=0$, where
$p^{(1)}_{E}=1+b+c-bc-d+bd+cd+bcd,$
$p^{(2)}_{E}=1-a-c-ac-e-ae-ce+ace,$
$p^{(3)}_{E}=1+a-b+ab+f-af+bf+abf,$
$p^{(4)}_{E}=1+d+e-de-f+df+ef+def,$
and $a,b,c,d,e,f\in\mathbb{T}.$
Proof.
Similar to Lemma 4.1 and Lemma 4.2, any constraint requiring orthogonality of the rows of $T_{8E}(a,b,c,d,e,f)$ can be factored into one of the equations above. ∎
As each of the equations contains only three of the six variables, the system is straightforward to solve.
Theorem 4.3.
Let $T_{8E}^{(3)}:=\mathcal{S}_{E}(\mathbb{T}^{3}),$ where
$\mathcal{S}_{E}(a,b,c)=T_{8E}(a,b,c,d^{\prime},e^{\prime},f^{\prime})$
with
$d^{\prime}=(-1-b-c+bc)/(-1+b+c+bc),$
$e^{\prime}=(-1+a+c+ac)/(-1-a-c+ac),$
$f^{\prime}=(-1-a+b-ab)/(1-a+b+ab).$
$T^{(3)}_{8E}$ is a 3-parameter family of complex Hadamard matrices.
Proof.
Consider the system of polynomial equations given in Lemma 4.3. Variables $d,\ e$ and $f$ can be solved with respect to $a,\ b$ and $c$ from the equations $p_{E}^{(1)}=0,\;p_{E}^{(2)}=0$ and $p_{E}^{(3)}=0$. The obtained solutions satisfy $p_{E}^{(4)}=0$ as well. By Lemma 2.2, $d^{\prime}$, $e^{\prime}$ and $f^{\prime}$ gets unimodular values whenever $a,b,c\in\mathbb{T}$. ∎
As the denominators of $d^{\prime}$, $e^{\prime}$ and $f^{\prime}$ have no roots in $\mathbb{T},$ the solution given in Theorem 4.3 contains all the solutions for the equations of Lemma 4.3.
4.4 Family $T^{(3)}_{8F}$
Consider a matrix-valued function $T_{8F}(a,b,c,d,e,f,g,h)=$
$\begin{bmatrix}1&ab&ef&abef&cd&dh&cg&gh\\ ab&1&-abef&-ef&dh&cd&-gh&-cg\\ ef&-abef&-1&ab&-cg&gh&cd&-dh\\ abef&-ef&ab&-1&-gh&cg&-dh&cd\\ be&-ae&bf&-af&1&-ch&dg&-cdgh\\ ae&-be&-af&bf&ch&-1&-cdgh&dg\\ bf&af&-be&-ae&-dg&-cdgh&1&ch\\ af&bf&ae&be&-cdgh&-dg&-ch&-1\\ \end{bmatrix}.$
Lemma 4.4.
$T_{8F}(a,b,c,d,e,f,g,h)$ is a complex Hadamard matrix if and only if if $p^{(1)}_{F}=0$, $p^{(2)}_{F}=0$, $p^{(3)}_{F}=0$, $p^{(4)}_{F}=0$, $p^{(5)}_{F}=0$ and $p^{(6)}_{F}=0$, where
$p^{(1)}_{F}=$ $-cd+b^{2}cd+be-bc^{2}e+bd^{2}e-bc^{2}d^{2}e-cde^{2}+b^{2}cde^{2},$
$p^{(2)}_{F}=$ $-ae+ad^{2}e+dh-a^{2}dh-de^{2}h+a^{2}de^{2}h+aeh^{2}-ad^{2}eh^{2},$
$p^{(3)}_{F}=$ $\ bf+bc^{2}f-cg-b^{2}cg+cf^{2}g+b^{2}cf^{2}g-bfg^{2}-bc^{2}fg^{2},$
$p^{(4)}_{F}=$ $-af-afg^{2}+gh+a^{2}gh+f^{2}gh+a^{2}f^{2}gh-afh^{2}-afg^{2}h^{2},$
$p^{(5)}_{F}=$ $\ abc^{2}d^{2}ef+abc^{2}efg^{2}-cdgh+a^{2}b^{2}cdgh-cde^{2}f^{2}gh$
$+a^{2}b^{2}cde^{2}f^{2}gh-abd^{2}efh^{2}-abefg^{2}h^{2},$
$p^{(6)}_{F}=$ $\ abef+abd^{2}efg^{2}+a^{2}cde^{2}gh-b^{2}cde^{2}gh+a^{2}cdf^{2}gh$
$-b^{2}cdf^{2}gh-abc^{2}efh^{2}-abc^{2}d^{2}efg^{2}h^{2},$
and $a,b,c,d,e,f,g,h\in\mathbb{T}$.
Proof.
All the orthogonality requirements can be factored to depend only on the given polynomials. ∎
We can notice that assignment $(a,b,c,d,e,f,g,h)=(a,b,b,e,e,f,f,a)$ satisfy the equations $p^{(1)}_{F}=0$, $p^{(2)}_{F}=0$, $p^{(3)}_{F}=0$ and $p^{(4)}_{F}=0$. Additionally, $-p^{(5)}_{F}(a,b,b,e,e,f,f,a)=p^{(6)}_{F}(a,b,b,e,e,f,f,a)=abef\cdot p_{B}(a^{2},e^{2},f^{2},b^{2})$. Thus, we are left with the orthogonality constraint $p_{B}(a^{2},e^{2},f^{2},b^{2})=0$, which can be solved in the same manner as in Theorem 3.3. Introducing suitable variable changes, we can notice that the matrices given by this solution are all equivalent to matrices in $T_{8B}^{(3)}$.
There are also matrices we have not seen before in $T_{8F}(\mathbb{T}^{8})$ and we need the following lemma to find them.
Lemma 4.5.
The system of equations given in Lemma 4.4 is equivalent to the system
[…]
where $a,b,c,d,e,f,g,h\in\mathbb{T}.$
Proof.
The polynomials $p^{(i)}_{F}$ are palindromic or anti-palindromic for all $i\in\{1,\ldots,6\}$. By Lemma 2.1, we can write $p_{F}^{(1)}=0$ as
$be-\frac{1}{be}+\frac{b}{e}-\frac{e}{b}+\frac{d}{c}-\frac{c}{d}+\frac{1}{cd}-cd=0$
This can be factored to
$\left(b-\frac{1}{b}\right)\left(e+\frac{1}{e}\right)-\left(c-\frac{1}{c}\right)\left(d+\frac{1}{d}\right)=0$
and by unimodularity of the variables, this is equivalent to
[…]
The other constraints can be dealt with in a similar manner. ∎
We can solve the system (4.1) by introducing different variables for real and imaginary parts and adding unimodularity constraints between them.
Theorem 4.4.
Let $T_{8F}:=\mathcal{S}_{F}(D)$, where
$\mathcal{S}_{F}(a_{r},b_{r},c_{i})=T_{8F}(a_{r}+ia_{i},b_{r}+ib_{i},c_{r}+ic_{i},d^{\prime},e^{\prime},f^{\prime},g^{\prime},h^{\prime}),$
with
$a_{i}=\sqrt{1-a_{r}^{2}},\;\;\;$ $b_{i}=\sqrt{1-b_{r}^{2}},$ $c_{r}=\sqrt{1-c_{i}^{2}},$
$d^{\prime}=d_{r}+id_{i},$ $d_{r}=\frac{b_{i}\sqrt{A_{2}}}{c_{i}\sqrt{B_{1}B_{2}}},$ $d_{i}=i\frac{a_{i}\sqrt{A_{1}B_{3}}}{|c_{i}|\sqrt{B_{1}B_{2}}},$
$e^{\prime}=e_{r}+ie_{i},$ $e_{r}=\frac{\sqrt{A_{2}}}{\sqrt{B_{1}B_{2}}},\quad$ $e_{i}=i\frac{a_{r}b_{i}\sqrt{A_{1}}}{\sqrt{B_{1}B_{2}}},$
$f^{\prime}=f_{r}+if_{i},$ $f_{r}=-\frac{a_{i}b_{r}|c_{i}|\sqrt{A_{1}}}{c_{i}\sqrt{B_{1}B_{4}}},\;\;$ $f_{i}=i\frac{|c_{i}|\sqrt{A_{2}}}{c_{i}\sqrt{B_{1}B_{4}}},$
$g^{\prime}=g_{r}+ig_{i},$ $g_{i}=\frac{a_{r}\sqrt{A_{1}B_{3}}}{c_{r}\sqrt{B_{1}B_{4}}},$ $g_{i}=i\frac{b_{r}|c_{i}|\sqrt{A_{2}}}{c_{r}c_{i}\sqrt{B_{1}B_{4}}},$
$h^{\prime}=h_{r}+ih_{i},$ $h_{r}=-\frac{a_{i}b_{r}c_{r}|c_{i}|}{\sqrt{B_{3}}c_{i}},$ $h_{i}=\frac{a_{r}b_{i}|c_{i}|}{\sqrt{B_{3}}},$
where
$A_{1}=(c_{i}-b_{i})(b_{i}+c_{i}),$ $A_{2}=-a_{i}^{4}b_{r}^{2}+a_{r}^{4}c_{i}^{2}+a_{i}^{2}b_{r}^{2}c_{i}^{2}-a_{r}^{2}b_{r}^{2}c_{i}^{2},$
$B_{1}=(a_{r}-b_{r})(a_{r}+b_{r}),$ $B_{2}=-a_{r}^{2}b_{r}^{2}+a_{r}^{2}c_{i}^{2}+c_{r}^{2},$
$B_{3}=a_{i}^{2}b_{r}^{2}+a_{r}^{2}c_{i}^{2}-b_{r}^{2}c_{i}^{2},\;\;$ $B_{4}=-a_{i}^{2}b_{r}^{2}-a_{r}^{2}c_{i}^{2},$
and
$D=\{(a_{r},b_{r},c_{i})\in[-1,1]^{3}\;|\;d_{r},e_{r},f_{r},g_{r},h_{r}\in[-1,1]\}.$
$T_{8C}^{(3)}$ is a 3-paremeter family of complex Hadamard matrices.
Proof.
One can confirm that the assignment given is a solution for the system of equations in Lemma 4.5. Unimodularity is guaranteed when all the real parts get values from $[-1,1]$ because $x_{r}^{2}+x_{i}^{2}=1$ holds for all of the variables. ∎
Any dephased matrix from the family $T_{8B}^{(3)}$ has at least six entries that are equal to $-1$ in its core. A typical dephased matrix from the family $T^{(3)}_{8F}$ has only two entries that are equal to $-1$. Thus, the solution presented in Theorem 4.4 gives us matrices inequivalent to the matrices of the previous solution that produced the family $T^{(3)}_{8B}$.
Once again, we find multiple solutions that differ only by signs; 128 in total. Numerical results suggest that all these solutions correspond to equivalent families. Similar considerations also indicate that the family $T^{(3)}_{8F}$ is likely symmetric.
5 Inequivalence
To show that the introduced families are pairwise inequivalent and not contained in any known families, we provide matrices that belong exclusively to these new families. Often, such considerations are done by utilising invariants such as defect [23], fingerprint [20] or Haagerup set [11]. We introduce an algorithm to determine the inclusion precisely. A more detailed study of such algorithms will be conducted in a separate study [17].
5.1 Algorithm for determining family membership.
In this section, we describe an algorithm that can be used to determine if an arbitrary matrix $H\in\mathbb{H}(n)$ has an equivalent matrix that belongs to a given family of complex Hadamard matrices. First, we introduce a condition that makes it straightforward to invert a matrix-valued function. We denote the set of the entries of a matrix $A=(a_{i,j})$ by $\{a_{i,j}\}$ and the multiset of entries by $\langle a_{i,j}\rangle$.
Let $f:\mathbb{T}^{k}\to M_{n}(\mathbb{C})$ be a matrix-valued function. We say $f(x_{1},\ldots,x_{k})$ is iteratively invertible if for all $i\in\{1,\ldots,k\}$ there exist an entry $\phi_{i}(x_{1},\ldots,x_{i})$ of $f$ such that fixing $x_{1},\ldots,x_{i-1}$ makes $\phi_{i}$ bijective in $x_{i}$. We refer to such entries $\phi_{1},\ldots,\phi_{k}$ as iterable entries of $f$.
Suppose that $f(x_{1},\ldots,x_{k})$ is an iteratively invertible function with iterable entries $\phi_{1},\ldots\phi_{k}$ at indices $(i_{1},j_{1}),\ldots,(i_{k},j_{k})$. We can now compute $(x_{1}^{\prime},\ldots,x^{\prime}_{k})$ from $H=f(x_{1}^{\prime},\ldots,x^{\prime}_{k})$ with
$(\phi^{-1}_{1}(h_{i_{1},j_{1}}),\;\phi^{-1}_{2}(\phi^{-1}_{1}(h_{i_{1},j_{1}}),h_{i_{2},j_{2}}),\ldots,\phi^{-1}_{k}(\phi_{1}^{-1}(h_{i_{1},j_{1}}),\ldots,h_{i_{k},j_{k}})),$ (5.1)
where $\phi_{i}^{-1}(x_{1},\ldots,x_{i})$ denotes the inverse map of $x_{i}\mapsto\phi_{i}(x_{1},\ldots,x_{i})$ for fixed values of $x_{1},\ldots,x_{i-1}.$
Let $H^{(k)}_{n}=f(D)$ be a family of complex Hadamard matrices, where $f:D\to\mathbb{H}_{d}(n)$ is an iteratively invertible function and $D\subseteq\mathbb{T}^{k}$. For a given matrix $H\in\mathbb{H}(n),$ we can decide if $H\in H^{(k)}_{n}$ with the following algorithm.
  1. Step 0.
    Choose iterable entries $\phi_{1},\ldots,\phi_{k}$ of $f$ and compute $\phi^{-1}_{1},\ldots,\phi^{-1}_{k}.$
  2. Step 1.
    Find all mutually permutation inequivalent dephased matrices $H^{\prime}$ equivalent to $H$.
  3. Step 2.
    For each $H^{\prime}=(h^{\prime}_{i,j})$, perform a backtracking search to find parameters $x^{\prime}=(x^{\prime}_{1},\dots,x^{\prime}_{k})$, such that $\langle f_{i,j}(x^{\prime})\rangle=\langle h^{\prime}_{i,j}\rangle$, by evaluating the formula (5.1) with elements from $\{h^{\prime}_{i,j}\}$.
  4. Step 3.
    For each pair $(H^{\prime},x^{\prime})$, determine if $f(x^{\prime}_{1},\ldots,x_{k}^{\prime})$ and $H^{\prime}$ are permutation equivalent.
If we find a permutation equivalence in Step 3 we can output permutation matrices $P_{1}$ and $P_{2}$ as well as an evaluation point $(x_{1}^{\prime}\ldots,x^{\prime}_{k})$ such that $\mathscr{D}(P_{1}HP_{2})=f(x_{1}^{\prime},\ldots,x_{k}^{\prime})$ implying that $H\in H^{(k)}_{n}$. If no permutation equivalence is found, then we have exhausted all possibilities and conclude that $H\notin H_{n}^{(k)}$. Permutation equivalence of two matrices can be determined with an algorithm from [17].
In practice, we need to do the arithmetic with floating-point numbers. Complex numbers $z_{1}$ and $z_{2}$ are considered equal if $|z_{2}-z_{1}|<\varepsilon$, where $\varepsilon$ is an accuracy parameter of the algorithm.
We do not need the functions $\phi_{i}$ to be necessary bijective in the whole domain $D$, but we can divide our considerations into different cases, usually concerning different branches of the square root. It is not difficult to find iterable entries $\phi_{1},\;\phi_{2},\;\phi_{3}$ in the functions $\mathcal{S}_{C},$ $\mathcal{S}_{D}$, $\mathcal{S}_{E}$ and $\mathcal{S}_{F}$. For example, for $\mathcal{S}_{E}$ we can choose the identity function three times given by the entries $a,\;b,$ and $c$. For $\mathcal{S}_{F}$ we can use for example functions $\phi_{1}(a)=[\mathscr{D}(\mathcal{S}_{F})]_{8,2}=1/a^{2}$, $\phi_{2}(b)=[\mathscr{D}(\mathcal{S}_{F})]_{7,2}=1/b^{2}$ and $\phi_{3}(a,b,c)=[\mathscr{D}(\mathcal{S}_{F})]_{5,3}=1/e^{\prime}(a,b,c)^{2}$ that are relatively easy to invert.
Example 5.1.
Consider an affine family $F_{4}^{(1)}:=F_{4}(\mathbb{T})$ given by
$F_{4}(a)=\begin{bmatrix}1&1&1&1\\ 1&ia&-1&-ia\\ 1&-1&1&-1\\ 1&-ia&-1&ia\\ \end{bmatrix}.$
Suppose that we want to determine whether a complex Hadamard matrix
$G=\begin{bmatrix}\gamma&\beta\gamma&\beta&-1\\ \overline{\beta}\gamma&\gamma&-1&\overline{\beta}\\ \overline{\beta}&-1&\gamma&\overline{\beta}\gamma\\ -1&\beta&\beta\gamma&\gamma\\ \end{bmatrix},$
where $\beta=\exp(\sqrt{-3})$ and $\gamma=\exp(\sqrt{-7})$, belongs to $F^{(1)}_{4}.$
The function $F_{4}(a)$ has only one parameter, so we only need to choose one iterable entry. We can choose $\phi_{1}(a)=[F_{4}(a)]_{2,2}=ia$ and obtain its inverse $\phi_{1}^{-1}(h)=-ih.$
In Step 1, we find all the dephased matrices that are equivalent to $G$ up to permutation equivalence. There are only two such matrices:
$G_{1}=\begin{bmatrix}1&1&1&1\\ 1&1&-1&-1\\ 1&-1&\gamma^{2}&-\gamma^{2}\\ 1&-1&-\gamma^{2}&\gamma^{2}\\ \end{bmatrix}\text{ and }G_{2}=\begin{bmatrix}1&1&1&1\\ 1&1&-1&-1\\ 1&-1&\gamma^{-2}&-\gamma^{-2}\\ 1&-1&-\gamma^{-2}&\gamma^{-2}\\ \end{bmatrix}.$
In Step 2, we check if there exists $h\in\{1,-1,\gamma^{2},-\gamma^{2}\}$ such that the multiset of the entries of $F_{4}(\phi_{1}^{-1}(h))$ equals the multiset of the entries of $G_{1}$. We find two such matrices $F_{4}(-i\gamma^{2})$ and $F_{4}(i\gamma^{2}).$ Similarly, for $G_{2}$ we find the matrices $F_{4}(-i\gamma^{-2})$ and $F_{4}(i\gamma^{-2}).$
Finally, in Step 3, we compute that $F_{4}(-i\gamma^{2})$ and $F_{4}(i\gamma^{2})$ are both permutation equivalent to $G_{1}$. Also, $F_{4}(-i\gamma^{-2})$ and $F_{4}(i\gamma^{-2})$ are both permutation equivalent to $G_{2}$. We can output any of the found equalities
$\mathscr{D}(P_{1}GP_{2})=F_{4}(i\gamma^{2}),\qquad\mathscr{D}(P_{2}GP_{2})=F_{4}(-i\gamma^{2})$
$\mathscr{D}(P_{3}GP_{2})=F_{4}(i\gamma^{-2})\;\text{ or }\mathscr{D}(P_{4}GP_{2})=F_{4}(-i\gamma^{-2}),$
where $P_{1}=(e_{1}\;e_{3}\;e_{4}\;e_{2})$, $P_{2}=(e_{1}\;e_{3}\;e_{2}\;e_{4})$, $P_{3}=(e_{4}\;e_{2}\;e_{1}\;e_{3})$, and $P_{4}=(e_{2}\;e_{4}\;e_{1}\;e_{3}),$ that all prove the inclusion $G\in F^{(1)}_{4}.$
When we run the same test with
$G^{\prime}=\begin{bmatrix}\gamma&\beta\gamma&\beta&-1\\ \overline{\beta}\gamma&\gamma&-1&\overline{\beta}\\ \overline{\beta}&-1&\gamma&\overline{\beta}\gamma\\ -1&\beta&-\beta\gamma&-\gamma\\ \end{bmatrix},$
which is not a complex Hadamard matrix, the algorithm proceeds the same way until Step 3. At that point, no satisfying permutations are found, and the algorithm returns a negative result.
In an implementation of the algorithm, a depth-first traversal is more natural, as it allows for immediate termination upon finding a valid solution.
5.2 Inequivalence of the families
The following result shows that we have indeed found previously unknown matrices.
Theorem 5.1.
Let $H^{(k)}_{8}$ be a $k-$parameter family of complex Hadamard matrices of order 8 from the list
$\begin{split}F^{(5)}_{8},\;D^{(5)}_{8A},\;D^{(5)}_{8B},\;&S_{8A}^{(4)},\;S^{(4)}_{8B},\\ T_{8B}^{(3)},\;T_{8C}^{(3)},\;T^{(3)}_{8D},\;(T^{(3)}_{8D})^{T},&\;T^{(3)}_{8E},\;(T^{(3)}_{8E})^{T},\;T^{(3)}_{8F}.\end{split}$ (5.2)
Let $f:D\to\mathbb{H}(8)$ be the function defining $H^{(k)}_{8}$ as given in Sections 3 and 4. Consider $H^{\prime}=f(e^{2i},e^{3i},\dots,e^{(k+1)i})\in\mathbb{H}(8)$. The only family from the list (5.2) to which $H^{\prime}$ belongs is $H_{8}^{(k)}$ itself.
Proof.
One can check that the point $(e^{2i},e^{3i},\dots,e^{(k+1)i})$ belongs to the domains of all the families in the list (5.2). An Exhaustive computer search with the algorithm described in Section 5.1 yields the result. ∎
An immediate corollary of Theorem 5.1 is that all the families in the list (5.2) are pairwise inequivalent, and no family is contained within another. One can also confirm that the point $(e^{2i},e^{3i},\ldots,e^{ki})$ is not a boundary point of any of the domains, meaning there exists a neighbourhood around that point belonging uniquely to the corresponding family.
5.3 Butson-type matrices
The classification of complex Hadamard matrices of Butson-type into equivalence classes has been completed for many small parameter values of $n$ and $q$. These results are extensively documented in [13]. A study of the Butson matrices suggests that there are still order 8 families to discover.
In [21], all BH$(8,4)$ Butson-type matrices are placed in some affine family from the list (3.1). For BH$(8,6)$ matrices, such a categorisation is not possible. Consider a BH$(8,6)$ matrix
$B_{1}=\begin{bmatrix}\omega^{0}&\omega^{0}&\omega^{0}&\omega^{0}&\omega^{0}&\omega^{0}&\omega^{0}&\omega^{0}\\ \omega^{0}&\omega^{0}&\omega^{0}&\omega^{0}&\omega^{3}&\omega^{3}&\omega^{3}&\omega^{3}\\ \omega^{0}&\omega^{1}&\omega^{3}&\omega^{4}&\omega^{0}&\omega^{2}&\omega^{3}&\omega^{5}\\ \omega^{0}&\omega^{2}&\omega^{5}&\omega^{3}&\omega^{4}&\omega^{5}&\omega^{2}&\omega^{1}\\ \omega^{0}&\omega^{3}&\omega^{2}&\omega^{5}&\omega^{2}&\omega^{5}&\omega^{1}&\omega^{4}\\ \omega^{0}&\omega^{3}&\omega^{4}&\omega^{1}&\omega^{5}&\omega^{2}&\omega^{0}&\omega^{3}\\ \omega^{0}&\omega^{4}&\omega^{1}&\omega^{3}&\omega^{3}&\omega^{2}&\omega^{5}&\omega^{0}\\ \omega^{0}&\omega^{5}&\omega^{3}&\omega^{2}&\omega^{1}&\omega^{5}&\omega^{4}&\omega^{2}\\ \end{bmatrix}$
and its inequivalent transpose $B_{2}=B_{1}^{T}$, where $\omega=\exp(2\pi i/8).$
Proposition 5.1.
The matrices $B_{1}$ and $B_{2}$ do not belong in any of the families in the list (5.2).
Proof.
This is a result of an exhaustive computer search with the algorithm described in Section 5.1. ∎
We can also find Butson matrices with $n=8$ and $q>6$ that do not belong to any of the known families of complex Hadamard matrices.
Acknowledgements
I would like to thank Professor Patric Östergård for his insightful guidance and valuable comments.